Question:


 For the data given below;

Years Price of Apples

 (in $) Quantity of Apples

 (in $) Price of Bananas (in $) Quantity of Bananas 

(in $) Prices of Mangoes (in $) Quantity of Mangoes 

(in $)

2010 4 80 5 100 10 200

2011 5 100 6 120 11 220

2012 6 120 7 140 14 280

2013 7 140 8 160 16 320


Find out:

Cost of CPI Basket 

Consumer Price Indices (CPI) for each year considering the year 2010 as base year

Inflation rate for each year

What do you mean by demand pull and cost push inflation. 


1
Expert's answer
2021-11-23T11:16:15-0500


From the table above:

P.A is the Price of Apples

Q.A is the Quantity of Apples

P.B is the Price of Bananas.

Q.B is the  Quantity of Bananas

P.M is the Prices of Mangoes

Q.M is the Quantity of Mangoes .

Cost of CPI Basket for each year is calculated as: "=P.A \times Q.A + P.B \times Q.B + P.M\times Q.M"

Consumer Price Indices (CPI) for each year is calculated as: "\frac{Cost\ of\ CPI\ basket\ for\ the\ year}{Cost\ of\ CPI\ basket\ for\ base\ year} \times 100"

Inflation rate for each year: "\frac{later\ CPI-earlier\ CPI}{earler\ CPI} \times 100"

Cost-push inflation is the decrease in the aggregate supply of goods and services stemming from an increase in the cost of production. Demand-pull inflation is the increase in aggregate demand, categorized by the four sections of the macroeconomy